OpenStudy (anonymous):

if tanA+tanB=x and cotA+cotB=y then prove that, cot(A+B)=(1/x - 1/y)

hartnn (hartnn):

hi, have you attempted it by yourself first ? and know the formula for tan(A+B) or cot (A+B ) ?

OpenStudy (dumbcow):

\[\cot(A+B) = \frac{1-\tan A \tan B}{\tan A + \tan B} = \frac{1}{x} - \frac{\frac{1}{\cot A \cot B}}{\frac{1}{\cot A} +\frac{1}{\cot B}}\] can you take it from there

hartnn (hartnn):

if you're comfortable with tan (A+B) formula, then write cot A + cot B = y in terms of tan, to get x/(tan A tan B) = y now you have tan A tan B and tan A + tan B in terms of x and y just rearrange